OpenStudy (anonymous):

3 = (6/F) * x rearrange for F can u please show all workings?

OpenStudy (anonymous):

pretty straight forward, im just hitting a mental block, please help

OpenStudy (anonymous):

Do you any options or choices with you? :)

OpenStudy (anonymous):

na, just need it rearranged.

OpenStudy (anonymous):

3 = (6/F) * x 0=(6/F)* x-3 F= (6x-18) I think not sure :)

OpenStudy (anonymous):

Very good, you are very very clever, thankyou E Cosbia, have I spoken to you before? anyway thanks for the help.

OpenStudy (anonymous):

u look sort of familiar, at least your name does, any way thanks for the help. Johnathon

OpenStudy (anonymous):

thanks for the help, u r truly a blessing, all the best.

OpenStudy (anonymous):

im sorry, what is your problem? i don't like to argue on these things im here to study.

OpenStudy (anonymous):

Your welcome :)